### v1

Abstract
Deep learning (DL) techniques have achieved significant success in various software engineering tasks (e.g., code completion by Copilot). However, DL systems are prone to bugs from many sources, including training data. Existing literature suggests that bugs in training data are highly prevalent, but little research has focused on understanding their impacts on the models used in software engineering tasks. In this paper, we address this research gap through a comprehensive empirical investigation focused on three types of data prevalent in software engineering tasks: code-based, text-based, and metric-based. Using state-of-the-art baselines, we compare the models trained on clean datasets with those trained on datasets with quality issues and without proper preprocessing. By analysing the gradients, weights, and biases from neural networks under training, we identify the symptoms of data quality and preprocessing issues. Our analysis reveals that quality issues in code data cause biased learning and gradient instability, whereas problems in text data lead to overfitting and poor generalisation of models. On the other hand, quality issues in metric data result in exploding gradients and model overfitting, and inadequate preprocessing exacerbates these effects across all three data types. Finally, we demonstrate the validity and generalizability of our findings using six new datasets. Our research provides a better understanding of the impact and symptoms of data bugs in software engineering datasets. Practitioners and researchers can leverage these findings to develop better monitoring systems and data-cleaning methods to help detect and resolve data bugs in deep learning systems.
1 Introduction
Deep Learning (DL) solutions have been widely adopted in many applications, including speech recognition amodei2016deep , software testing mao2019extensive ; cummins2018compiler , autonomous driving grigorescu2020survey , and software development yang2022dlinse . However, deep learning models face unique challenges due to their complex architectures and data-driven paradigms arpteg2018software . Compared to traditional logic-driven software, DL-based software follows a data-driven computing paradigm where its models are trained using data. The training data can be noisy, biased, or incomplete, which may lead to unexpected or erroneous behaviours in the DL models whang2020data . Besides, deep learning models can be very complex and opaque, making it difficult to understand their decision process. Thus, the poor quality of training data in the DL models could pose significant challenges to their reliability and trustworthiness.
Debugging deep learning systems is a challenging task. DL bugs can originate from various sources, such as training data, hyperparameters, and model structure, and can lead to system crashes and unexpected runtime behaviour islam2019fse . They can also result in severe consequences, as evidenced by the fatal accidents involving self-driving cars selfdrivingcarcrash ; teslaAutopilotCrash . The non-determinism of DL systems leads to different outcomes across multiple runs, making debugging challenging nagarajan2018impact . According to existing literature islam2019fse , DL bugs can be divided into five categories: Data, Model, Structural, Non-Model Structural and API Bug.
Among these categories, data bugs have been reported as the most prevalent ones, accounting for 26% of all bugs in deep learning systems islam2019fse . They originate from errors in training data, such as incorrect labels, duplicates, out-of-distribution records, and missing values islam2019fse ; taxonomyRealFaults ; cote2024data . These errors can significantly affect the model performance liang2022advances ; sambasivan2021everyone . Resolving data bugs is very challenging yin2023dynamic ; wang2024empirical . Data bugs are hidden in the dataset and implicitly affect a model’s behaviour. They also propagate to the model parameters during training, which makes their detection difficult. Furthermore, according to existing work croft2022data , many existing benchmark datasets constructed by human annotators contain up to 70% mislabeled data, which leads to data bugs in the DL models relying on those datasets croft2022data .
Existing studies tantithamthavorn2015impact ; mahmood2015impact ; kim2011dealing ; wu2021data ; xu2023data ; croft2023data have investigated the impact of data quality issues (e.g., label noise, class imbalance) on model performance. However, they primarily focus on the quality of the raw training data and overlook two crucial aspects: the preprocessing stage and the model training process. First, data preprocessing is essential for preparing the data in a suitable format for training. Errors, biases, or loss of information during preprocessing can propagate through the whole development steps of a model, significantly degrading the model’s performance amato2023data . Despite its importance, the impact of preprocessing on model performance has not been thoroughly investigated by the existing literature. Second, training is a crucial step in model development. Observing the training process and monitoring a model’s internal state, such as gradients, weights, and biases, can offer valuable insights into how data quality issues affect the model’s learning process. For example, if the model is experiencing vanishing or exploding gradients, they might indicate poor feature scaling or outliers in the data glorot2010understanding . Similarly, analyzing the weights and biases of the model during training can indicate whether the model is learning meaningful patterns or is being misled by noisy or corrupted data points koh2017understanding . Such insights can help practitioners take appropriate actions to mitigate the impact of data bugs in their DL-based applications. However, no existing work examines the impact of data bugs on the training behaviours of DL models. Our study aims to address this critical gap in the literature.
In this paper, we conduct an empirical study to investigate the impact of data quality and preprocessing issues on the training of deep learning models used in software engineering tasks. First, we select three types of data for our analysis based on their frequent use in software engineering tasks: code-based, text-based, and metric-based. Second, we select state-of-the-art baseline models using these data types and compare their faulty models (containing data bugs) with corresponding bug-free versions. Third, we capture the detailed training logs using the Weights & Biases framework and analyze the training metrics and model properties (e.g., gradients, weights, and biases) to derive meaningful insights. We perform qualitative and quantitative analyses of the gradients, weights, and biases to identify symptoms and manifestations of data quality and preprocessing issues. Finally, we validate our findings with new datasets to ensure the generalizability of our results. Thus, we answer four important research questions as follows:
RQ1: How do data quality and preprocessing issues in code-based data affect the training behavior of deep learning models?
We investigate the impact of data quality issues from two code-based datasets, Devign and BigVul zhou2019devign ; fan2020ac . These datasets contain C/C++ functions from multiple projects and have known quality issues croft2023data . Our analysis reveals that these issues cause gradient instability during training and significantly impair a model’s capacity to capture complex patterns. Furthermore, our manual analysis of attention weights shows that data quality issues in the code-based data can reduce the code comprehension abilities of the deep learning model.
RQ2: How do data quality and preprocessing issues in text-based data affect the training behaviour of deep learning models?
We assess the impact of data quality issues from two text-based datasets, Eclipse, maintained in Bugzilla and Hadoop, maintained in JIRA. Our study reveals that quality issues in text data lead to reduced abnormal weight distributions and overfitting of models to noisy patterns during model training. Our qualitative analysis using the t-SNE plots highlights the deep learning models’ difficulty learning consistent feature representations from noisy data.
RQ3: How do data quality and preprocessing issues in metric-based data affect the training behavior of deep learning models?
We investigate the impact of data quality issues from two metric-based datasets, Openstack and QT, which frequently suffer from class imbalance problems mcintosh2018fix ; hoang2019deepjit . Our analysis reveals that quality issues in metric-based data lead to vanishing gradients and poor optimization during the training process. Additionally, we perform qualitative analysis using GradCAM, an explainable AI technique for visual analysis of model inputs. Our qualitative analysis demonstrates that models trained on imbalanced data focus on irrelevant tokens and struggle to generalize to unseen data.
RQ4: How well do our findings on data quality and preprocessing issues generalize to other code-based, text-based, and metric-based datasets?
We evaluate the generalizability of our findings using six new datasets: D2A and Juliet (code-based), Spark and Mozilla (text-based), and Go and JDT (metric-based). Our analysis reveals that the data quality and processing issues in the training data lead to reduced model capacity, gradient instability, overfitting, and biased learning of the models, which align with our above findings. In contrast, the models trained on cleaned datasets show none of these issues. Such observations increase confidence in our findings and underscore the challenges of data bugs in deep learning models used in software engineering tasks.
Paper Organization: The remainder of the paper is organized as follows. Section 2 provides background knowledge about deep learning and explainable AI. Section 3 describes the methodology of our study in detail, including data type selection, data collection, experimental setup, post-hoc analysis and quantitative analysis. Section 4 talks about the findings for the different research questions. Section 5 discusses the implications of our findings for various stakeholders. Section 6 reviews the related literature. Finally, Section 7 presents threats to the validity of our study, and Section 8 concludes the paper.
2 Background
2.1 Data Quality Issues in Software Engineering Datasets
Label Noise: Label noise refers to errors in the labels assigned to data instances in a dataset. These errors can stem from various sources, including insufficient information, annotator mistakes, subjective judgments, and data encoding issues frenay2014comprehensive . Label noise is prevalent in real-world datasets and thus can significantly impair the DL models using those datasets frenay2013classification . This issue is also relevant in software engineering datasets since they are often used to train DL models supporting code search, vulnerability detection, and program understanding wang2024empirical ; xu2024code ; nie2023understanding . For example, in a dataset for vulnerability detection, non-vulnerable code could be mistakenly labelled as vulnerable or vulnerable code could be labelled as non-vulnerable. Such label errors often result in inaccurate, biased, or flawed models nie2023understanding . In this work, we train multiple deep learning models using datasets containing label noise to determine its impact on the models.
Class Imbalance: Class imbalance refers to a disproportionate representation of different classes within a dataset. It is also highly prevalent in various software engineering datasets, including API recommendation, code review automation, and defect prediction irsan2023multi ; tufano2022using ; song2018comprehensive . For example, in defect prediction, there are often significantly fewer defective modules than non-defective ones. Existing studies mcintosh2018fix ; kamei2012large ; keshavarz2022apachejit ; mahbub2023defectors show that only 5%-26% of the files contain the defective instances. When trained on imbalanced datasets, DL models tend to be biased towards the majority class and demonstrate poor performance in identifying the minority class giray2023use .
Data Obsolescence: Data obsolescence, also known as concept drift, refers to the evolution of data over time and is a significant challenge for software engineering datasets. As software systems evolve, the characteristics of their data change, which leads to concept drift. This phenomenon is prevalent in many datasets, including the ones used for log-level prediction, anomaly detection and duplicate bug report detection ma2018robust ; ouatiti2024impact ; zhang2023duplicate . Recent research by Zhang et al. zhang2023duplicate revealed that most duplicate bug report detection techniques were only evaluated using data up to January 2014. When these same techniques were applied to more recent data, their performance decreased significantly, highlighting the impact of data obsolescence on the effectiveness of deep learning models.
2.2 Model Explanation
Attention-Based Analysis: Attention mechanisms help deep learning models focus on the most relevant parts of the input vaswani2017attention . Their inherent ability to assign importance weights to different input elements can also be leveraged to interpret the behaviours of the models. Recently, attention weights of input have been used to investigate the explainability of the deep learning models solving software engineering tasks fu2022linevul ; pornprasit2022deeplinedp ; fu2022gpt2sp . For example, Fu et al. fu2022linevul leveraged the self-attention mechanism to explain the predictions of their proposed technique for vulnerability detection. In this study, we employ their dataset and attention-based analysis to examine how data quality and preprocessing issues affect a DL model’s training behaviour and learning capacity.
t-SNE: t-Distributed Stochastic Neighbor Embedding (t-SNE) van2008visualizing is a technique that can reduce high-dimensional feature representations learned by neural networks and visualize them in 2D or 3D plots. These plots are often used to inspect how well a model has learned to differentiate among different classes in the data. By comparing t-SNE visualizations for different models and examining their separation of classes, we can evaluate their capacities to learn feature representations for a given task. This technique has also been used in several software engineering tasks, including duplicate bug report detection messaoud2022duplicate . As duplicate bug report detection is one of our selected tasks, we utilise t-SNE to explain and visualize the predictions of our trained model for this task.
Grad-CAM: Grad-CAM (Class Activation Mapping) is a technique that visualizes the input features contributing the most to a neural network’s outputs selvaraju2017grad . By analyzing the feature weights in the final convolutional layer of a convolutional neural network, Grad-CAM produces a heatmap highlighting the important regions in the input relevant to the model’s output. In our study, we employ DeepJIT hoang2019deepjit as a baseline technique for defect prediction. Since DeepJIT is a CNN-based technique, we use Grad-CAM to understand and visualize the impact of input features on the model output.
3 Methodology
Fig. 1 shows the schematic diagram of our empirical study. We discuss different steps of our study as follows.
3.1 Data Type Selection
Selecting different types of data is crucial for our study since each data type has unique characteristics and quality issues. By examining multiple types of data, we can comprehensively investigate how data bugs manifest and affect deep learning models. Based on the prevalence in software engineering datasets, Yang et al. yang2022dlinse identified the three most prevalent types of data: code-based, text-based, and metric-based. These data types have unique characteristics as follows.
Code-based data: Code-based data is frequently used in training deep learning models that target various software engineering tasks, such as code clone detection, code generation, program repair, and vulnerability detection fu2022linevul ; zhang2020learning ; li2017cclearner ; tian2020evaluating . This data type encompasses source code files, test cases, and code changes yang2022dlinse .
Text-based data: Natural language texts play a crucial role in numerous software engineering tasks. Existing studies have used text-based data in deep learning techniques for software engineering tasks yang2022dlinse . Text-based data includes requirements specifications, design documents, code comments, commit messages, bug reports, user reviews, and question-answer posts from forums like Stack Overflow he2020duplicate ; ouatiti2024impact ; mondal2024can .
Metric-based data: Metric-based data comprises various statistics derived by static analysis tools (e.g., SonarQube, Understand) from various software repositories. They quantify different aspects of source code, software design, development process, and software quality. Metric-based datasets have been used in several software engineering tasks such as defect prediction, effort estimation, code smell detection, and software maintainability assessment hoang2019deepjit ; liu2019deep ; choetkiertikul2018deep .
Thus, based on the prevalence of data types and their relevance to software engineering tasks, we consider code-based, text-based and metric-based data for our study (Step 1, Fig. 1).
3.2 Study Design
3.2.1 Task Selection
Selecting representative tasks for each type of data above is a critical step in our analysis (Step 2, Fig. 1). By carefully choosing the tasks that use our selected data types, we can examine how data bugs manifest and affect the training of deep learning models. In particular, we have selected the following representative tasks for our analysis, as suggested by Yang et al. yang2022dlinse :
(a). Vulnerability Detection: Detection of vulnerabilities in software code is one of the key applications where deep learning models show promising results. These models often use source code or binary for their detection task.
(b). Duplicate Bug Report Detection: Duplicate bug report detection is a prominent use case of text-based data from software engineering that is leveraged by DL models. The task’s objective is to identify bug reports that describe the same underlying issue from previous bug reports.
(c). Defect Prediction: Predicting defects in code components using software metrics is a common application of deep learning in software engineering. This task involves predicting the defect proneness of software modules based on various code metrics (e.g., complexity, coupling, cohesion).
Figure 1: Schematic diagram of our study
3.2.2 Baseline Selection
After selecting the representative tasks, we choose multiple baseline approaches for the selected tasks. These approaches provide baseline deep learning models, which are used in our subsequent analysis. To select baselines for each of the types of data, we used the following criteria: (a) the technique should use a neural network model to perform one of the three tasks above, and (b) the technique should provide a comprehensive replication package with source code, datasets, and instructions for reproducibility. Furthermore, we prioritized the baselines with clean and buggy versions in their replication packages. Based on these filtration criteria, we selected the following state-of-the-art baselines for our study.
(a) Code-based data: LineVul fu2022linevul for vulnerability detection. LineVul is a state-of-the-art transformer-based model for line-level vulnerability detection in code. The model takes code tokens as input and learns to predict the vulnerability status of each line.
(b) Text-based data: DCCNN he2020duplicate for duplicate bug report detection. DCCNN, short for Duplicate bug report detection using Convolutional Neural Networks, employs a CNN-based architecture to identify duplicate bug reports from their textual descriptions. The CNN model learns to extract relevant features and patterns from a pair of bug reports and then predict whether they are duplicates or not.
(c) Metric-based data: DeepJIT hoang2019deepjit for defect prediction. DeepJIT is an end-to-end deep learning framework for Just-In-Time (JIT) defect prediction. This technique computes software metrics from datasets of code changes and extracts salient features from commit messages. The learned features from commit messages and code changes are encoded into numerical matrices and then processed by separate CNN layers to predict whether the commit will likely introduce a defect.
Besides the relevance to the selected data types, these baselines have demonstrated state-of-the-art performance in their respective tasks, which justifies their selection for our study (Step 3, Fig. 1).
3.2.3 Datasets
We utilized six diverse datasets that were used by deep learning models targeting our tasks above: Devign and BigVul (code-based), Eclipse and Hadoop (text-based), and OpenStack and QT (metric-based). These datasets vary significantly in size, composition, and specific characteristics, allowing for comprehensive training and evaluation of deep learning models across different tasks. For each type of data, we chose two datasets to perform data source triangulation, per the guidelines by Runeson et al. runeson2009guidelines , and to have multiple sources of evidence as recommended by Yin et al. yin2009case . Table. 1 provides a detailed summary of each dataset, including their sizes and brief descriptions, offering a clear overview of the data used in our experiments.
Table 1: Summary of datasets used in our study
Data Type Dataset Size Description
Code-Based Devign zhou2019devign 27,318 functions Vulnerable functions from 5 open-source projects in C/C++
BigVul fan2020ac 188,636 functions Vulnerable code snippets from 211 projects in multiple languages
Text-Based Eclipse lazar2014generating 74,376 bug reports Bug reports from Eclipse project (2001-2007)
Hadoop hadoop 14,016 bug reports Bug reports from Hadoop project (2006-2013)
Metric-Based OpenStack mcintosh2018fix 66,065 source files 21 code metrics per file from OpenStack project
QT mcintosh2018fix 95,758 source files 21 code metrics per file from QT project
3.3 Experimental Setup
This section describes our experimental methodology, including how we configure our system, implement the baselines, and observe training behaviours.
3.3.1 System Configuration
To reflect the original environments of our selected baselines, we use the following setup:
(a) Code Editors: We use Visual Studio Code v1.79.0, which is a popular code editor for building DL-based applications ide .
(b) Dependencies: To automatically detect the API libraries used in the baseline techniques, we use the pipreqs package  [^1] https://pypi.org/project/pipreqs/. We also install the dependencies for each baseline into a separate virtual environment using the venv  [^2] https://docs.python.org/3/library/venv.html module.
(c) Frameworks: We use Tensorflow and PyTorch for our experiments, as DeepJIT hoang2019deepjit and LineVul were originally developed in PyTorch, whereas DCCNN was originally developed in Tensorflow.
(d) Python Version: For our experiments, we use the same versions of Python originally used by the authors when they published their work.
(e) Hardware Config: Our experiments were run on a Compute Canada node having a Linux (CentOS 7) Operating System with 64GB primary memory (i.e., RAM) and 16GB GPU Memory (NVIDIA V100 Tensor Core GPU).
3.3.2 Baseline Preparation
To prepare the baselines, we utilized the replication packages provided by the authors of the original studies fu2022linevul ; hoang2019deepjit ; he2020duplicate . This allowed us to accurately reproduce the baseline setups and maintain the integrity of their original implementations. We prepared three variants of the original baselines for our analysis, which are described below.
1. Baseline with Clean Data: First, we obtained clean datasets from existing studies croft2023data ; zhang2023duplicate ; issta2021deepjit and prepared the original techniques to be trained using this high-quality data (Step 4, Fig. 1). This variant represents the ideal setup for the models as it is prepared with clean and high-quality data.
2. Baseline with Buggy Data: To investigate the impact of data quality on model performance, we created a second variant prepared with data quality issues. We obtained buggy datasets from the same existing studies croft2023data ; zhang2023duplicate ; issta2021deepjit and configured the baseline models to be trained using this buggy data (Step 5, Fig. 1). This step will allow us to observe how bugs in the training data might affect the model’s learning process and subsequent performance.
3. Baseline with Missing Preprocessing: For our third variant, we focused on preprocessing faults, which are among the most prevalent types of faults in deep learning, according to the existing taxonomy taxonomyRealFaults . We removed the preprocessing operations from the training pipeline in preparation for training (Step 6, Fig. 1). By setting up the model to simulate the missing preprocessing operation, we aim to understand how the absence of these crucial steps might impact a model’s performance and robustness.
3.3.3 Integrating Weights and Biases
To better understand the training process and model behaviour, we integrated Weights and Biases (W&B) logging wandb into the existing baselines and their variants (Step 7, Fig. 1). After successfully integrating W&B, we trained the models and monitored and captured their higher-level training, such as the model’s loss and accuracy during the training. Additionally, we logged lower-level statistics such as gradients, weights, and biases from all model layers. This allowed us to track a model’s progress and performance over time. Furthermore, we logged the runtime configurations to capture the computational environment in which the experiments were conducted. Finally, to account for the variability and stochasticity inherent in neural networks, we ran each baseline model and corresponding variants fifteen times arcuri2014hitchhiker .
3.4 Quantitative Analysis
To quantitatively analyze the impact of data quality and preprocessing issues on the deep learning models, we leveraged the data from Weights and Biases (W&B) captured during the training process. We use the gradients, weights, and biases over multiple runs and calculate aggregated statistics for each layer of the models (Step 7, Fig. 1). Table. 2 discusses the aggregate metrics used for analysis. By examining these aggregated statistics, we identified symptoms and manifestations of data quality and preprocessing issues and quantified their impact on deep learning models. We analyzed the statistical differences between models trained on clean datasets and those trained on buggy datasets to identify if there were any differences in their performance. We observed how data bugs and preprocessing issues affected different layers’ weights, biases, and gradients. By quantifying the impact on specific parts of the models, such as attention layers, convolutional layers, and fully connected layers, we gained insights into how data quality and preprocessing issues can degrade the performance and reliability of deep learning models in software engineering tasks.
3.5 Post-Hoc Analysis
To gain deeper insights into the impact of data bugs on our deep learning models, we conducted a post-hoc analysis using explainable AI techniques (Step 8, Fig. 1). These techniques help us understand how the models make predictions and identify the influential parts of the input data.
For the code-based data, we analyzed LineVul’s attention weights. Since the LineVul model is attention-based, an analysis of its attention weights helps us identify the parts of the input code that the model focuses on when making predictions about vulnerabilities. By comparing the attention weights from models trained on clean and buggy data, we observed how data bugs affect a model’s attention and potentially lead to incorrect predictions.
For the text-based data, we utilized t-SNE (t-Distributed Stochastic Neighbor Embedding) visualization van2008visualizing . t-SNE is a dimensionality reduction technique that helps us visualize high-dimensional data in a lower-dimensional space. By applying t-SNE to the representations learned by the DCCNN model at different layers, we visualized the decision boundary for duplicate and non-duplicate bug reports. This helps us understand how the learned representations differ between models trained on clean and buggy data, providing insights into the impact of data bugs on the model’s learning process.
For the metric-based data, we implemented GradCAM selvaraju2017grad . In the context of defect prediction using DeepJIT, GradCAM helps us visualize which metrics and parts of the input features are most influential in the model’s decision. By contrasting the GradCAM outputs of models trained on clean data against those trained on buggy data, we identify how data bugs affect the model’s focus and potentially lead to incorrect defect predictions.
Table 2: Aggregate Metrics for Model Analysis
Operator Description
max The maximum value of the property in the layer.
min The minimum value of the property in the layer.
median The median value of the property in the layer.
mean The average value of the property in the layer.
var The variance of the property in the layer.
std The standard deviation of the property in the layer.
skew A measure of the asymmetry of the distribution of the property in the layer.
kurt A measure of the peakedness and tail heaviness of the property’s distribution in the layer.
spar The fraction of properties in a layer that are zero or close to zero.
3.6 Validating the Derived Findings
To assess the generalizability of our findings from RQ1, RQ2, and RQ3, we conducted additional experiments using six new datasets not included in our initial analysis (Step 9, Fig. 1), following similar approaches in existing studieshossain2024deep . In particular, we chose six new datasets representing three types of data: two code-based datasets (Juliet and D2A), two text-based datasets (Mozilla and Spark), and two metric-based datasets (Go and JDT). We collect the buggy and clean versions of these datasets from the same studies zhang2023duplicate ; issta2021deepjit ; croft2023data , which were used in our previous steps. Table  3 summarises our validation datasets, including their sizes, compositions, and brief descriptions.
Table 3: Summary of datasets used for validation and generalization
Data Type Dataset Size Description
Code-Based Juliet 253,002 test cases C/C++ and Java test cases covering 181 CWEs
D2A 1,295,623 samples Samples from six open-source projects such as OpenSSL, FFmpeg etc.
Text-Based Mozilla 193,587 bug reports Bug reports from Mozilla projects
Spark 9,579 bug reports Bug reports from Apache Spark project
Metric-Based Go 61,224 files Metrics from Go programming language project
JDT 13,348 files Metrics from Eclipse Java Development Tools
In our validation, we retrained each of the baseline models (LineVul, DCCNN, and DeepJIT) with their corresponding datasets containing clean and buggy data. We performed similar quantitative analyses, including monitoring training behaviour and examining model components. Moreover, we also perform the qualitative analyses using the same post-hoc techniques (attention weight analysis, t-SNE visualization, and GradCAM) as in our original study. By comparing the results from these validation experiments with our initial findings, we aimed to determine whether the observed impacts of data quality issues are consistent across different datasets within each data type, and this assessed the generalizability of our conclusions.
4 Study Findings
In our study, we examined the effects of four common data bugs: label noise, class imbalance, concept drift, and missing preprocessing. To address each research question, we constructed and trained a comprehensive dataset of 120 buggy models, with 30 models dedicated to each of the four bug types. Additionally, we trained 30 models on clean data to serve as a baseline, helping us establish the expected training behaviour. We used the W&B logging framework wandb to capture the training behaviours of both faulty and clean models. This section presents our findings, focusing on the most common symptoms of these data bugs across three data types: code-based, text-based, and metric-based.
4.1 RQ1: How do data quality and preprocessing issues in code-based data affect the training behaviour of deep learning models?
4.1.1 Impact of Data Quality
(a) Inconsistent Learning: Inconsistent learning in neural networks refers to the phenomenon where different layers or components of a network learn at varying rates or effectiveness, leading to suboptimal overall performance bengio1994learning . We observe inconsistent learning across the neural network layers when models were trained on low-quality data. As shown in Table. 4, 80% of the models trained on data with label noise and 56.67% of the models trained on data with concept drift demonstrated inconsistent learning. Our manual analysis of these models shows that $\approx$20of their layers were affected, especially the attention and output layers. Due to the data bugs, these layers learn very slowly, as shown by their near-zero biases ($<$ 0.01) neal2018modern . In contrast, the layers unaffected by data quality issues learn quickly (bias >=>=> = 0.5). Attention and the output layers of the models also showed normal bias values (bias >=>=> = 0.5). Thus, the disparity in the bias values across the different layers of a neural network trained on low-quality code data results in inconsistent learning.
(b) Reduced Model Capacity: We found that models trained on low-quality data exhibited a reduced capacity to capture complex patterns from data. As shown in Table 4, 86.67% of models trained on data with label noise and 63.33% of models trained with concept drift demonstrated reduced capacity. Our analysis of these models revealed that 90-95% of their layers were affected, particularly embedding, attention, and dense layers. These layers exhibited notably small weights (e.g., $\mu\approx 0.011$, $\mu\approx 0.047$ for models trained with label noise and concept drift, respectively). These values were significantly lower than the typical weight values of 0.1 to 0.3, which were observed when the models were trained on clean data. The small and tightly clustered weights suggest a limited range in the representational power of a model’s neurons, which indicates the reduced capacity of the corresponding model liu-etal-2020-understanding ; bengio1994learning .
(c) Gradient Instability: We also noted gradient instability when models were trained on low-quality data. As shown in Table 4, 73.33% of models trained on data with label noise and 53.33% with class imbalance exhibited gradient instability. Our analysis of these models revealed that 30-40% of their layers were affected, especially the attention and dense layers. These layers demonstrated extreme gradient values, ranging from $1\times 10^{-9}$ to $1\times 10^{1}$, which was a significant deviation from the range of $1\times 10^{-3}$ to $1\times 10^{-1}$, as observed in the baseline models trained on clean data. This wide range of gradient magnitudes leads to unstable learning, where some layers update their weights rapidly, and others remain almost static. This gradient instability results in suboptimal model performance, as the model fails to develop a coherent understanding of code data across all of its layers pascanu2013difficulty .
Table 4: Manifestations of Data Bugs in Code-Based Models
Data Quality Issues Incosistent Learning Reduced Model Capacity Gradient Instability
Label Noise 80.00% 86.67% 73.33%
Class Imbalance 6.67% 23.33% 53.33%
Concept Drift 56.67% 63.33% 16.67%
4.1.2 Impact of Missing Preprocessing Operations
(a) Slow Convergence: In code-based data, preprocessing operations such as tokenization and stopword removal are crucial, as they enable the accurate parsing and analysis of source code. Our analysis reveals that 66.67% of the models trained with missing preprocessing operations had slower convergence. We also found that $\approx$15of the layers in these models suffered from slow convergence, especially the output layers. The slow convergence of the parameters of the output layers indicates that no class was probable enough, which affects the decision-making ability of a model. Their weights and gradients were close to zero ($<$ $10^{-6}$) as opposed to $1\times 10^{-3}$ to $1\times 10^{-1}$ from the baseline models burden1997numerical . Such low gradients and weights often result in inefficient gradient flow, which affects a model’s ability to converge towards the optimal parameter values.
(b) Incorrect Learning: We noted that the lack of preprocessing operations in code-based data led to inconsistent or incorrect learning. Our analysis showed that the models trained on clean datasets exhibited a uniform and bounded distribution in their bias values (e.g., ranging from 0.34 to 0.72). Such a distribution enables better learning and generalization to unseen data srivastava2014dropout . On the contrary, we noted unbounded bias values in the models trained on data without preprocessing. Some layers displayed bias values near zero ($\approx$0), while others had bias values exceeding 1.0. This behavior was prevalent in 53.33% of the models trained without preprocessing, affecting $\approx$60of their layers. These inconsistent bias patterns indicate that the models struggle to leverage biases effectively, resulting in a limited set of learned weights primarily clustered near the origin ($\approx$0). Consequently, this leads to poor generalization and reduced “guessing power”, degrading the model’s ability to comprehend and analyze source code effectively Bennett_2017 .
(c) Distorted Distributions: Missing preprocessing steps can significantly distort the distribution of weights from deep learning models, as suggested by our analysis. We found that 70% of the models trained without preprocessing were affected. Our analysis of these models showed that 80% of their layers exhibit distorted distributions, particularly in embedding, attention, and dense layers. They demonstrated a skewness of $>1$ and a kurtosis of $|k|>>>3$, which are significantly higher than the skewness (0.4) and kurtosis (1.3) of the model weights trained on clean data. Such high skewness and extreme kurtosis indicate that the weights and gradients of an affected model are concentrated at a point farther right from the baseline counterparts. That is, small changes in input can cause a large change in output, which makes the model highly sensitive to input changes, reducing its robustness and reliability.
4.1.3 Impact analysis of Label Noise using Attention Mechanisms
Attention weights have been effectively used in software engineering research to understand the correlation between the input and a model’s predictions (outputs) cao2024systematic . In our study, we also analyze the attention weights of a model to understand the impact of data quality issues on the model’s training. Our study examined various data quality issues, and we found that label noise had the strongest impact on models trained with code-based data (see Table 4). Based on this finding, we decided to conduct a qualitative analysis to gain a deeper understanding of how label noise affects code-based data. To investigate this issue, we trained 60 models in total. We divided these models into four groups of 15 models each: (a) models trained on a clean Devign dataset, (b) models trained on a clean BigVul dataset, (c) models trained on a Devign dataset with label noise, and (d) models trained on a BigVul dataset with label noise. For each group, we calculated the average attention weights across all 15 models. This approach helps us identify consistent patterns and mitigate the impact of non-determinism in deep learning. We discuss our findings below.
Bug-free Data: Models trained on bug-free datasets on vulnerability detection demonstrate a strong focus on tokens critical for identifying potential security flaws. This focus is evident in the BigVul dataset, which contains source code from Chrome, Linux, Android, and Tcpdump. When analyzing models trained on this dataset, we observe high attention weights on key functions and data structures related to security vulnerabilities. For instance, rds6_inc_info_copy (87.3796) and struct rds_incoming (64.1523) receive significant attention, indicating their relevance to network-related vulnerabilities. The models also heavily emphasize network fields such as laddr (85.7431) and faddr (86.2221), demonstrating their awareness of potential security issues in network communications. The Devign dataset contains source code from Linux, FFmpeg, Qemu, and Wireshark. Models trained on this dataset focus on memory-related operations and type conversions, which are one of the primary causes of security vulnerabilities Stevens_Fenner_Rudoff_2013 . Tokens like uint16_t (88.8645) and uint64_t (86.9365) receive very high attention weights, which reflects their importance to proper data handling. Critical function calls for cross-platform compatibility, such as le16_to_cpu (82.2068) and le64_to_cpu (80.8885), are also strongly emphasized. Models trained on both datasets show significant attention to error-checking patterns and memory operations like sizeof (30.5034) and if(!conn) (32.1079). These attention weights suggest the models have developed a comprehensive understanding of code elements related to security flaws from the datasets.
Buggy Data: In contrast, the models trained on buggy data showed their inability to focus on tokens crucial for vulnerability detection. In BigVul, the model assigns significant weights to non-vulnerable tokens or variables such as int (37.6882), rt (36.5832), and nl (41.7406). Similarly, in Devign, the model focuses its attention on various elements like static (50.7911), CCIDBus (55.7003), and Device (43.4922). In other words, the model fails to prioritize the tokens that might signal potential vulnerabilities, such as unsafe operations or improper error handling. Moreover, the models trained on data with label noise pay considerable attention to non-informative tokens, such as newline characters (26.1465 in BigVul, 54.1878 in Devign) and empty tokens (8.2194 in BigVul, 10.1193 in Devign). This diffused attention pattern, evident across both datasets, suggests that these models struggle to distinguish between benign code patterns and those that may introduce vulnerabilities. This limits their effectiveness in identifying security-critical code sections.
Summary of RQ1: Models trained on code-based datasets containing data bugs exhibit many issues, including inconsistent learning across layers, reduced model capacity and gradient instability. Furthermore, missing preprocessing can lead to slow convergence, incorrect learning patterns, and distorted distributions, impairing the models’ ability to understand and process code effectively.
4.2 RQ2: How do data quality and preprocessing issues in text-based data affect the training behaviour of deep learning models?
4.2.1 Impact of Data Quality Issues
(a) Abnormal Weight Distribution: We observed abnormal layer weights in the models trained on low-quality text data. As shown in Table 5, 93.33% of the models trained on data with concept drift exhibited this phenomenon. Our analysis of these models revealed that 80-85% of their layers were affected, particularly the convolutional and dense layers. These layers demonstrated unusually high weight variances (e.g., $\sigma^{2}\approx 2.3$), significantly deviating from their baseline counterparts (e.g., mean weight variance $\sigma^{2}\approx 0.64$). The observed abnormal weight distribution can be attributed to the dynamic nature of text data in software engineering. As software and technology evolve, the corresponding text-based data also changes, which causes shifts in the underlying data distribution. This evolution in software and text can affect how the text is represented, and the relevant weights and biases are learned by the models. As a consequence, older and newer data points might require different sets of optimal parameters. During model training, the text data with concept drift generate large gradient signals to push the parameter values towards the optimal settings for their respective time periods, which could have led to abnormal weight distribution in the models.
(b) Gradient Skewness: We also observed significant gradient skewness in the models trained on low-quality text data. As shown in Table 5, 76.67% of the models trained on data with concept drift and 56.67% with label noise exhibited this phenomenon. Our analysis revealed that 80-90% of the layers in these models were affected, particularly the convolutional layers and batch normalization layers. The gradient distribution of these layers demonstrated substantial skewness ($\gamma\approx 2.9$ for models trained with concept drift, and $\gamma\approx 1.2$ for models trained with label noise), far exceeding the baseline skewness values ($\gamma\approx 0.2$). As discussed above, concept drift in text-based data can lead to large gradient signals during training. This not only affects weight distribution but also could make the gradients skewed. Large gradient signals shift the overall distribution towards higher values, causing positive skewness. Similarly, label noise can also cause gradient skewness. In our datasets with label noise, similar data points have different labels due to incorrect labelling. When a model encounters these mislabeled examples during training, it tries to learn a decision boundary that separates these similar points with different labels, which is theoretically impossible. This leads to large prediction errors, as the model’s output significantly differs from the (incorrectly assigned) target labels. These large errors, in turn, result in large gradient updates during backpropagation. As the training process continues, these inconsistent and large gradient updates accumulate, causing the overall gradient distribution to be skewed.
(c) Overfitting: We noticed overfitting in the models trained on low-quality text data. As shown in Table 5, 86.66% of models trained on data with class imbalance exhibited overfitting. Our analysis revealed that 60-65% of the layers in these models were affected, particularly the dense layers. The median bias value of these layers ($med\approx 1.1$) is much higher than their baseline counterparts ($med\approx 0.07$). When a dataset has a disproportionate number of samples in one class compared to the others, the model tends to become biased towards the majority class. As a result, the model’s biases become skewed towards the majority class, which results in the model overfitting the training data.
Table 5: Manifestations of Data Bugs in Text-Based Models
Data Quality Issues Abnormal Weight Distribution Gradient Skewness Overfitting
Label Noise 16.67% 56.67% 3.33%
Class Imbalance 36.67% 23.33% 86.66%
Concept Drift 93.33% 76.67% 43.66%
4.2.2 Impact of Missing Preprocessing Operations
(a) Numerical Instability: Models trained on raw text data showed clear signs of numerical instability. According to our analysis, 63.33% of the models exhibited this issue. From our analysis, we observed that 80-90% of their layers were affected, particularly in the embedding layers and convolutional layers. These layers showed extreme weight ranges (-2.16 to 2.54), significantly deviating from the typical ranges of -0.1 to 0.1 for weights observed in the models trained on the clean dataset. These extreme weight parameter values indicate potential instability in the models and can be traced to the lack of proper text preprocessing and tokenization brownlee_2017 . Without these steps, the model encounters a high frequency of out-of-vocabulary words, including typos, rare technical terms, and inconsistent word forms (e.g., different tenses or misspellings of the same word). As the model attempts to learn representations for these out-of-vocabulary words, it often resorts to extreme weight adjustments, particularly in the embedding and convolutional layers, which are responsible for feature extraction. This leads to high volatility and significant variance in the layer weights, which results in exploding gradients and numerical instability in the neural networks bengio1994learning ; goodfellow2016deep .
(b) Skewed Bias Distributions: Models trained on raw texts demonstrated significant skewness in their bias distributions. Our analysis revealed that 53.33% of the models exhibited this issue, with 50-55% of the layers affected, particularly the convolutional layers. These layers demonstrated high skewness ($>$ 1) and moderate kurtosis ($|k|>$ 3) in their bias values, which are significantly higher than their baseline counterparts (skewness = 0.34, kurtosis = 1.82), collected from the models trained on clean, preprocessed data. This phenomenon can be attributed to the imbalanced representation of the text data due to a lack of preprocessing operations. Without tokenization and stop word removal, certain stop words or phrases may appear disproportionately in the dataset. This imbalance directly affects the bias values in the model. As certain words or phrases occur more frequently, the neurons associated with these common patterns are activated more often. Consequently, the biases for these neurons receive more frequent updates during training, pushing them further from their initial values. In contrast, biases for neurons associated with less common words or patterns receive fewer updates. This disparity in update frequency leads to a spread in the bias values, with some shifting significantly and others remaining closer to their starting points. This results in a skewed distribution of bias values, which reflects the imbalanced nature of the input data.
4.2.3 Qualitative Analysis: Impact of Concept Drift on Decision Boundary
To assess the impact of data quality issues on deep learning models’ training behaviour, we also performed a qualitative analysis using t-SNE plots. First, we trained 30 models on clean datasets and 30 models on datasets containing concept drift. We selected concept drift for our qualitative analysis, as it affected the highest number of models (see Table. 5). Then, we generated the t-SNE plots for the dense layer representations, as shown in Fig. 2 and Fig. 3. We focus on dense layers since their decision boundaries capture complex, non-linear relationships between features and directly contribute to the final classification. Furthermore, existing literature messaoud2022duplicate has analyzed the t-SNE plots of dense layers in their qualitative analysis. We discuss the outcome of our analysis below.
From Fig. 2 and 3, we note the impact of concept drift on the dense layers of our analyzed deep learning models for duplicate bug report detection. In the first dense layer (Fig. 2(a)), models trained on clean data show a clear separation between the two classes, indicating effective capture of the patterns. On the other hand, models trained on buggy data (Fig. 3(a)) exhibit complex representations with ambiguous decision boundaries. This observation is supported by an analysis of KL Divergence, which serves as the cost function of the t-SNE plot. As the cost function, KL Divergence is automatically calculated and optimized during the generation of these visualizations, with lower KL Divergence values indicating a better representation of the high-dimensional data in the 2D space. For the first dense layer, the mean KL divergence for buggy models (0.32) is higher than that of bug-free models (0.26), indicating more noise in the layer representation of models affected by concept drift. The second dense layer (Fig.2(b) and Fig.3(b)) demonstrates similar trends, with bug-free models showing well-defined decision boundaries and buggy data models displaying unclear decision boundaries and fragmented clusters. This trend is further corroborated by the KL divergence analysis, where buggy models exhibit a mean KL divergence of 0.24 compared to 0.14 for bug-free models in the second dense layer. The increase in KL divergence for buggy models across both layers (23% in the first layer and 71% in the second layer) underscores the concept drift’s impact on the network. It also highlights how concept drift progressively affects a model’s ability to learn meaningful representations and make accurate predictions, ultimately impairing its performance in duplicate bug report detection.
(a) First dense layer
(b) Second dense layer
Figure 2: t-SNE plots for models trained on bug-free data
(a) First dense layer
(b) Second dense layer
Figure 3: t-SNE plots for models trained on buggy data
Summary of RQ2: Quality issues in text-based data introduce many challenges for deep learning models, including abnormal weight distribution, gradient skewness, and overfitting. Moreover, missing preprocessing techniques can lead to numerical instability and skewed bias distributions, worsening a models’ ability to learn effectively from the bug reports.
4.3 RQ3: How do data quality and preprocessing issues in metric-based data affect the training behaviour of deep learning models?
Table 6: Manifestations of Data Bugs in Metric-Based Models
Data Quality Issues Sparse Parameter Updates Vanishing Gradients Poor Optimization
Label Noise 33.33% 36.67% 23.33%
Class Imbalance 76.67% 83.33% 66.67%
Concept Drift 13.33% 16.67% 10.00%
4.3.1 Impact of Data Quality
(a) Sparse Parameter Updates: We observed sparse updates to the parameters of neural networks trained on low-quality metric data. As shown in Table 6, 76.67% of models trained with class imbalance exhibited this phenomenon. Our analysis revealed that 50-55% of their layers were affected, particularly the fully connected layers. In these layers, up to 85% of the neurons showed no change in their gradients. That is, 85% of the neurons in a particular layer have not received any updates and have effectively stopped learning. This phenomenon can be explained by the disproportionate representation of classes in imbalanced datasets. In such datasets, the model is exposed to significantly more examples from the majority classes, causing neurons associated with these classes to receive frequent and substantial gradient updates. Conversely, neurons detecting minority classes receive updates far less frequently and with smaller magnitudes. As a consequence, a significant portion of neurons, particularly those responsible for minority class features, become ‘dead’ and do not contribute to the learning process. Thus, the model struggles to learn representations for minority classes, leading to poor performance in these underrepresented categories.
(b) Vanishing Gradients: We also observed the problem of vanishing gradients in models trained on imbalanced datasets. As shown in Table 6, 83.33% of the models trained on imbalanced datasets exhibited this phenomenon. In these models, gradient magnitudes decreased exponentially from the output layer towards the input layer. Similarly, models trained on corrupted data experienced gradient drops by a factor of $10^{6}$ between the output and input layers. In contrast, models trained on clean data showed a moderate drop in gradients, with only a $10^{2}$ factor decrease from output to input. As discussed above, class imbalance leads to sparse parameter updates, which in turn contributes to the vanishing gradient problem. As neurons associated with minority classes receive infrequent and smaller updates, their gradient signals become progressively weaker as they propagate backwards from the output layer through the network. Consequently, the early layers of the network (e.g., input layers), crucial for learning fundamental features, receive negligible updates for minority class patterns and suffer from vanishing gradients. This phenomenon severely impairs a model’s ability to learn from minority classes, further exacerbating the issues caused by class imbalance.
(c) Poor Optimization: We also noticed poor optimization in the models trained on low-quality metric data. As shown in Table 6, 66.67% of models trained on datasets with class imbalance suffered from this issue. Our analysis revealed that the models trained on low-quality data led to a stagnation of loss values after the first epoch. This contrasts sharply with models trained on balanced datasets, which exhibited steady and consistent decreases in loss throughout training. The average training and validation loss for models trained on imbalanced datasets was approximately 1.5 times higher than those trained on clean datasets. This trend was consistent across different software projects. For instance, in the Openstack project, the training loss value increased from $\approx$75 for clean data to $\approx$115 for buggy data. Similarly, in the QT project, the training loss value rose from $\approx$110 for clean data to $\approx$170 for buggy data. When a dataset has a class imbalance, the model’s corresponding optimization process is severely affected. The stagnation in loss values, coupled with the vanishing gradient problem, suggests that the model might get stuck in local minima. As a result, the model struggles to learn representations for minority classes or refine its decision boundaries, leading to poor performance.
4.3.2 Impact of Missing Preprocessing Operations
(a) Exploding Gradients: Models trained on raw numerical data showed signs of exploding gradients. According to our analysis, 66.67% of the models suffered from this issue. We observed that 70-80% of their layers were affected, particularly in the output layers of the network. These layers showed extreme gradient values (e.g., from -4.23 to 2.76), significantly deviating from the typical gradient ranges of -1 to 1 observed in the models trained on properly scaled datasets. These extreme gradient values can be explained by the lack of appropriate feature scaling and normalization glorot2010understanding . Without these preprocessing steps, the model encounters features with vastly different scales, leading to disproportionate weight updates. This leads to high volatility and significant variance in the gradient magnitudes, which results in exploding gradients and potential divergence from the normal behaviours in the neural networks pascanu2013difficulty ; goodfellow2016deep .
(b) High Variance in Weight Distribution: Models trained on unprocessed numerical data demonstrated significant variance in weight distribution. Our analysis revealed that 53.33% of the models exhibited this issue, with 85-90% of their layers affected, particularly the input and early hidden layers. These layers demonstrated high variance in weight magnitudes ($\sigma^{2}\approx 5.76$), which is significantly higher than their baseline counterparts ($\sigma^{2}\approx 0.89$) collected from the models trained on normalized data. This high variance in weight distribution stems from the interaction within a neural network and the diverse scales inherent in software metrics. For example, lines of code might be in thousands, while cyclomatic complexity might be in single digits. During training, the neural network compensates for these scale disparities to ensure a fair representation of all features in the learning process lecun2002efficient . It assigns larger weights to small-scale features (e.g., cyclomatic complexity) to amplify their effect and smaller weights to large-scale features (e.g., lines of code) to prevent them from dominating the output. This compensation mechanism results in a wide range of weight values. The input and early hidden layers are particularly susceptible to this issue as they directly interact with the raw feature values lecun2002efficient , contributing to the observed weight variance issue.
4.3.3 Qualitative Analysis: Impact of Class Imbalance on Feature Importance
To analyze the impact of class imbalance on the buggy and bug-free models, we apply the GradCAM (Gradient-weighted Class Activation Mapping) technique selvaraju2017grad . This approach has been used in previous studies to understand the effectiveness and efficacy of defect prediction models issta2021deepjit . For our GradCAM analysis, we used datasets specifically designed for defect prediction, focusing on software metrics. Our dataset exhibited an imbalanced distribution, with the minor class comprising non-defective instances and the major class consisting of defective instances. By examining the GradCAM outputs, we investigate which metrics the models prioritize and how this focus differs between the two types of models trained on two distinct codebases: Openstack and QT.
Buggy Models: Our analysis of the GradCAM output for the buggy models reveals a disproportionate focus on obscure or overly specific tokens across both codebases. In the OpenStack model, tokens like ‘audit_location_generator’ (0.9987) and ‘i62ce43a330d7ae94eda4’ (0.9962) have probabilities very close to 1, while more general programming concepts like ‘if’ (0.0023) and ‘return’ (0.0041) have probabilities near 0. Similarly, the Qt model exhibits high probabilities for tokens such as ‘Q_OBJECT_FAKE’ (0.9976) and ‘qAsConst’ (0.9953), while potentially bug-indicative tokens like ‘nullptr’ (0.0037) or ‘delete’ (0.0052) have very low probabilities. This pattern suggests that the severe class imbalance has likely led to overfitting, causing the models to associate rare, noise tokens, or codebase-specific terms with software defects in the training data. Consequently, these buggy models may fail to generalize and capture the broader semantics and context of software defects across different codebases.
Bug-free Models: In contrast, according to GradCAM output, the bug-free models focus on a wider variety of relevant technical terms and phrases in both OpenStack and Qt codebases. For OpenStack, tokens such as ‘vendor-data’ (0.9871), ‘xenapinfs-glance-integration’ (0.9923), ‘live_migr-ate’ (0.9905), ‘deprecationwarning’ (0.9889), ‘regressions’ (0.9917), and ‘backward’ (0.9894) have probabilities close to 1. These terms intuitively relate to areas where bugs occurred in OpenStack code, such as migrations, deprecations, regressions, and vendor integrations. Similarly, for Qt, we observe a focus on both framework-specific and general programming concepts. Tokens like ‘QObject’ (0.9934), ‘connect’ (0.9912), ‘emit’ (0.9901), ‘virtual’ (0.9887), and ‘override’ (0.9923) have probabilities close to 1. Additionally, memory-related terms such as ‘new’ (0.9876), ‘delete’ (0.9908), and ‘shared_ptr’ (0.9892) also show high probabilities, reflecting the importance of memory management in C++ code. This observation suggests that the bug-free models learn to focus on both codebase-specific concepts and general programming patterns that might have historically been associated with software defects. By mitigating the data quality issues, particularly class imbalance, these models have developed more generalizable representations that meaningfully capture the semantics of software defects across different codebases and frameworks.
The stark contrast in token focus between the buggy and bug-free models highlights the significant impact of class imbalance on model behavior. While the buggy models emphasize seemingly arbitrary or overly specific tokens to learn about software defects, the bug-free models focus on relevant technical concepts and terms.
Summary of RQ3: The presence of data quality issues in metric-based data introduces several challenges for deep learning models, including sparse parameter updates, vanishing gradients, and poor optimization. Moreover, inadequate preprocessing techniques can lead to exploding gradients and high variance in weight distribution, further hindering the models’ ability to learn from the data effectively.
4.4 RQ4: How well do our findings on data quality and preprocessing issues generalize to other code-based, text-based, and metric-based datasets?
To evaluate the generalizability of our findings, we conducted experiments on six additional datasets: D2A and Juliet (code-based), Spark and Mozilla (text-based), and Go and JDT (metric-based). We present our findings below.
In code-based datasets, we found that 83.33% of models (25 out of 30) exhibited inconsistent learning, 86.67% (26 out of 30) showed reduced model capacity, and 76.67% (23 out of 30) experienced gradient instability when trained on data with label noise or concept drift. For text-based datasets, 93.33% of models (28 out of 30) showed abnormal weight distribution, 76.67% (23 out of 30) demonstrated gradient skewness, and 86.67% (26 out of 30) exhibited overfitting when trained on data with concept drift or class imbalance. In metric-based datasets, 76.67% of models (23 out of 30) showed sparse parameter updates, 83.33% (26 out of 30) experienced vanishing gradients, and 73.33% (22 out of 30) demonstrated poor optimization when trained on data with label noise or class imbalance. When models were trained without appropriate preprocessing of data, we also observed similar patterns across all dataset types. In code-based datasets, 66.67% of models (20 out of 30) trained without proper preprocessing showed slow convergence, and 56.67% (17 out of 30) exhibited incorrect learning. For text-based data, 66.67% of models (20 out of 30) trained on unprocessed data showed numerical instability, while 56.67% (17 out of 30) exhibited skewed bias distributions. In metric-based datasets, 73.33% of models (22 out of 30) trained without preprocessing demonstrated exploding gradients, and 56.67% (17 out of 30) showed high variance in weight distribution. All percentages observed in these new datasets align closely with our original findings, varying within a 1-5% margin. These consistencies in observations across diverse datasets underscore the generalizability of our findings, highlighting the persistent impact of data quality issues and the importance of proper preprocessing in model training for software engineering tasks.
4.4.1 Model’s Behaviour on Clean Datasets
After addressing data quality issues and retraining the above models, we observed significant improvements in model behaviour and performance across all types of datasets: code-based, text-based, and metric-based. The impacts of data bugs were significantly diminished, and the models exhibited stable and normal behaviours during their training.
In code-based datasets, we found only 3.33% of models (1 out of 30) showing signs of inconsistent learning, compared to 83.37% previously. Model capacity improved substantially, with only 6.67% of models (2 out of 30) failing to capture complex patterns. Gradient stability also increased, with only 10.00% of models (3 out of 30) demonstrating instable gradients, down from 76.67% in the presence of data quality issues.
For text-based datasets, the chance of models having abnormal weight distribution decreased dramatically, with only 6.67% of models (2 out of 30) showing this behavior, compared to 93.33% before. Gradient skewness was largely prevented, with only 10.00% of models (3 out of 30) demonstrating skewed gradient distributions. Overfitting was significantly reduced, with only 6.67% of models (2 out of 30) showing this issue, down from 86.67% previously.
In metric-based datasets, sparse parameter updates were nearly prevented, with only 6.67% of models (2 out of 30) showing any signs of this issue. The vanishing gradient problem was significantly mitigated, with only 10.00% of models (3 out of 30) demonstrating this issue. Poor optimization was largely addressed, with only 13.33% of models (4 out of 30) showing this problem, compared to 73.33% previously.
The preprocessing issues, once addressed, also led to substantial improvements across all dataset types. In code-based datasets, only 6.67% of models (2 out of 30) showed slow convergence or incorrect learning patterns. For text-based data, only 10.00% of models (3 out of 30) demonstrated numerical instability or skewed bias distributions. In metric-based datasets, only 13.33% of models (4 out of 30) had exploding gradients or high variance in weight distribution.
4.5 Statistical Signifiance Tests
To assess the generalizability of our findings, we performed statistical significance tests with the following hypotheses.
[…]
symptoms are independent
[…]
symptoms are dependent
We selected McNemar’s test pembury2020effective for the statistical tests due to our experimental design and paired nominal data, as it specifically evaluates changes in binary outcomes (presence/absence of symptoms) within the same models. While McNemar’s test establishes statistical significance to test our null hypothesis that data quality issues and model symptoms are independent, the odds ratio measures the strength of their dependence szumilas2010explaining . In our analysis, an odds ratio of 1 would support our null hypothesis, which indicates that there is no association between data quality issues and model symptoms. Conversely, odds ratios greater than 1 would support our alternative hypothesis, suggesting that the presence of data quality issues increases the likelihood of observed symptoms. Moreover, odds ratios below 1 would indicate that data quality issues decrease the likelihood of model symptoms.
McNemar’s test results consistently rejected the null hypothesis across all symptoms ($p<0.01$). The odds ratios, measuring the strength of association between data quality issues and symptoms, were substantially greater than 1 in all cases, ranging from $15.53$ for abnormal weight distribution to $806.10$ for gradient skewness, strongly supporting our alternative hypothesis. The most pronounced dependencies were observed in inconsistent learning, gradient instability, and overfitting ($p<0.0001$), indicating that these symptoms were exclusively present when data quality issues existed. Even the weakest association, found in the symptom ”high variance in weight distribution” ($p=0.0098$), demonstrated a statistically significant association between the symptom and data quality issues.
These results demonstrate that addressing data quality issues and applying appropriate proper preprocessing operations to training data can significantly improve model training for various software engineering tasks. The consistent improvements observed across different dataset types further validate the generalizability of our findings and underscore the critical importance of data quality and preprocessing in machine learning applications for software engineering tasks.
Summary of RQ4: Our findings on data quality and preprocessing issues generalize consistently across diverse software engineering datasets, including D2A and Juliet (code-based), Spark and Mozilla (text-based), and Go and JDT (metric-based). The repeated emergence of similar patterns and enhancements after addressing the data quality issues across multiple datasets increases confidence in our study findings.
5 Implications for Stakeholders
The findings of our study have significant implications for various stakeholders involved in developing and deploying deep learning models targeting software engineering tasks. We discuss these implications in detail below.
Researchers: Researchers can reuse our dataset and leverage our empirical insights concerning label noise, concept drift, and class imbalance to develop robust data cleaning and preprocessing techniques. Our findings on the impact of missing preprocessing operations can also guide researchers in designing appropriate preprocessing pipelines tailored to different data types (code-based, text-based, and metric-based). Additionally, our analysis of how data quality issues affect the training of DL models can inform the development of more effective debugging techniques for DL models. Furthermore, researchers can integrate the explainable AI (XAI) techniques demonstrated in our study, such as attention visualization, t-SNE, and GradCAM, into their methodologies to better understand model behaviour and the impact of data quality issues. Finally, the researchers can use the symptoms of data bugs as verification criteria to enhance the reproducibility of deep learning bugs, as highlighted by existing literature shah2025towards .
Software Engineers: Software engineers can leverage our insights and implement continuous data monitoring systems that track properties such as weight distributions, gradient patterns, and learning stability across different layers. These systems can be used to automatically detect anomalies in training patterns and data distributions, enabling early identification of potential data quality issues before they are deployed in real-world systems. Our insights on the effects of concept drift can be used to establish processes for regular model retraining and deployment, ensuring their relevance over time. Moreover, the XAI techniques from our study can help software engineers debug and analyze model behaviour, identify potential biases, and address any issues before model deployment.
Deep Learning Engineers: Deep learning engineers can leverage our empirical findings to establish automated data quality assurance, design validation mechanisms and integrate them into MLOps pipelines. The identified data quality issues and their manifestations can serve as the basis for designing effective validation rules, statistical tests, and domain-specific constraints. Furthermore, our study emphasizes the importance of robust preprocessing pipelines, which can guide deep learning engineers in building modular and extensible preprocessing components that adapt to different data types and evolving data characteristics. Our findings enable deep learning engineers to develop optimized training procedures with built-in validation mechanisms to ensure stable model convergence. Additionally, they can implement comprehensive monitoring and alerting systems that track critical indicators like weight distributions, gradient flows, and learning patterns, allowing rapid detection and response to potential data quality issues during model training.
6 Related Work
Data quality is crucial for the performance and reliability of deep learning models in software engineering. Previous studies have investigated the effects of data quality issues like label noise frenay2013classification ; frenay2014comprehensive ; nie2023understanding ; croft2023data , class imbalance li2022robust ; liu2022comparative , data duplication zhang2023duplicate ; croft2023data ; lopez2024inter , and concept drift zhang2023duplicate ; lai2021towards ; schelter2015challenges on deep learning models.
Recent studies have also analyzed the impact of data quality issues on deep learning models for software engineering tasks. Wu et al. wu2021data investigated mislabeled instances in five publicly available datasets commonly used for security bug report prediction, including Chromium, Ambari, Camel, Derby, and Wicket, where they identified 749 security bug reports incorrectly labelled as non-security. They also reported significant performance improvements for a retrained model that was trained on correctly labelled data. Tantithamthavorn et al. tantithamthavorn2015impact analyzed over 3,900 issue reports from Apache projects and demonstrated that label noise can significantly impact the recall of a model. They also reported a significant improvement in the recall ($\approx$60%) for a retrained model on cleaned datasets. Kim et al. kim2011dealing evaluated the impact of intentionally injected noise into the datasets for defect prediction models through controlled experiments. They found that when $\approx$30of the data was mislabeled, the performance of defect prediction models decreased significantly, which highlights the models’ sensitivity to label noise. Fan et al. fanimpact investigated mislabeled changes in just-in-time defect prediction datasets and found that certain labelling approaches can lead to performance reduction of up to 5%. Xu et al. xu2023data leveraged adversarial learning to improve the data quality of obsolete comment datasets, which led to an improvement in the accuracy of multiple existing models by 18.1%. Croft et al. croft2023data conducted a systematic analysis of data quality in software vulnerability datasets and found out that 71% of the labels are incorrect and 17-99% of data points are duplicated across four state-of-the-art datasets. Furthermore, they also analyzed the impact of the data quality issues on vulnerability detection. They found that the model’s performance dropped by 65% when trained on clean data, which shows how the duplication and mislabelling of the data points can lead to inflated results.
In summary, the existing studies have examined the impact of data quality on model performance. However, a clear understanding of how these issues affect the training behaviours of a DL model remains limited. Most existing studies focus on specific data types or individual quality issues, preventing a comprehensive understanding of how data quality impacts deep learning models. Besides, how missing preprocessing operations can affect model behaviour is not well understood to date. Our study addresses these gaps in the literature through a systematic analysis of 900 models and 12 datasets targeting software engineering tasks.
Our work investigates how data quality and preprocessing issues affect the training behaviours of deep learning models. We adopt a comprehensive approach by examining three major issues - label noise, class imbalance, and concept drift - across software engineering data in code, text, and metric formats. We also demonstrate the generalizability of our findings through the reproduction of our findings on separate datasets, which were not used in our primary analysis. This multifaceted approach informs us of the impact of data bugs on deep learning models targeting software engineering tasks, addresses the gaps in the existing literature and encourages future efforts for appropriate debugging solutions.
7 Threats to Validity
One potential threat to the internal validity of our findings is the choice of baseline models and datasets. Although we followed systematic criteria for selection (Section 3.1, Section 3.2), there may be other relevant models or datasets not considered. We mitigate this threat by carefully documenting our selection criteria and validating our findings across diverse datasets within each data type category (code-based, text-based, and metric-based). Another threat is the potential for errors or biases in our analysis techniques. We mitigate this threat by using analysis techniques which are well-established in the literature and triangulating our findings through qualitative analysis (using XAI techniques) and quantitative analysis (using gradients, weights, and biases). Another threat is the subjective nature of our qualitative analysis, particularly in interpreting attention weights for code comprehension, t-SNE visualizations, and GradCAM results. We mitigated this by following established guidelines for analysis from previous studies issta2021deepjit ; fu2022linevul ; messaoud2022duplicate .
The main threat to external validity is the generalizability of our findings to software engineering tasks or data types that were not those considered. Our analysis is based on three tasks – vulnerability prediction, defect prediction and duplicate bug report detection. The impact of data quality and preprocessing issues may manifest differently in other tasks or application domains. Additionally, we focused on specific types of data quality issues, which may limit the generalizability of our findings. To mitigate these threats, we selected prevalent tasks, data types, and data quality issues based on comprehensive surveys and prior studies yang2022dlinse ; zhang2023duplicate ; croft2023data ; li2022robust ; liu2022comparative and used established benchmark datasets from different projects and domains.
The primary threat to conclusion validity comes from the inherent stochasticity in deep learning model behaviours. To account for this, we run each model multiple times and collectively analyze our findings in an aggregate manner across different data types and model behaviours, examining patterns in training dynamics (through gradients, weights, and biases). Our conclusions are drawn from these consistent, aggregate patterns observed in model behaviours when encountering data quality issues.
The primary threat to conclusion validity is the potential for statistical errors or violations of assumptions in our quantitative analysis. While we followed best practices and ran multiple trials to account for stochasticity, there may be inherent limitations or biases that could affect the accuracy or generalizability of our conclusions. Additionally, our conclusions are based on specific experiments and analysis techniques, and other approaches or methodologies could yield different or complementary insights. Regarding reproducibility, we acknowledge several potential threats. Environmental factors such as hardware specifications, software versions, and operating systems may influence results. To mitigate these threats, we provide detailed documentation of our experimental setup and a comprehensive replication package containing all source code, datasets, and configuration files used in our study replicationPackage .
8 Conclusion and Future Work
Deep learning systems suffer from data bugs that can significantly impact their reliability and trustworthiness. These bugs are particularly challenging to detect and resolve as they originate from various sources, including training data quality issues and preprocessing errors. In this study, we investigate the impact of data quality and preprocessing issues on the training behaviours of deep learning models targeting software engineering tasks. We target three prevalent data types: code-based, text-based, and metric-based data in the context of three tasks: vulnerability detection, duplicate bug report detection, and defect prediction. First, we replicated three state-of-the-art baselines, introducing data quality issues, omitting preprocessing operations and capturing their training-time behaviours using advanced tools (e.g., Weights & Biases). We then employed advanced explainable AI techniques to gain insights into the models’ learning processes and internal representations. Our key findings shed light on the impacts of data quality issues, such as label noise, concept drift, and class imbalance. These issues can lead to biased model behaviour, inconsistent learning across layers, reduced capacity to capture complex patterns, gradient instability during model training, overfitting to noisy patterns, vanishing gradients, poor optimization, and impaired generalization capabilities.
Additionally, the absence of necessary preprocessing operations can result in errors in parameter initialization, incorrect learning patterns, and biased feature extraction. This study emphasizes the critical importance of high-quality data and robust preprocessing pipelines for effective training of deep learning models in software engineering tasks. Our derived insights serve as a foundation for future work in developing data cleaning techniques, continuous data quality monitoring systems, and automated preprocessing pipelines tailored to different data types within the field of software engineering. Our findings can benefit debugging solutions by highlighting potential data-related issues that may impact model performance, developing more robust preprocessing techniques, and informing the design of automated quality assessment tools for machine learning pipelines in software engineering. These insights can also help practitioners identify and mitigate common data quality issues before negatively impacting model training and deployment.
Data Availability Statement (DAS)
All the data generated or analyzed during this study are available on Zenodo to help reproduce our results replicationPackage .
Conflict of Interest
The authors declare that they have no conflict of interest.